OpenStudy (anonymous):

Please help? A gambler has developed a weighted six-sided die with sides marked 1, 2, 3, 4, 5, 6. The probability of rolling a 6 is 0.4 and all other sides are equally likely. (X is a random variable). What is the probability that an even number is rolled?

OpenStudy (javalos01):

hello

OpenStudy (javalos01):

do you need help

OpenStudy (anonymous):

Yes, please

OpenStudy (anonymous):

I do not understand how to find this probability

OpenStudy (javalos01):

well do you know probabilities

OpenStudy (anonymous):

Yes I do, but what I am confused about is that he gave us an assigned probability. Usually the probability for rolling a dice is 1/6, but he made it .4 and that confuses me.

OpenStudy (javalos01):

ok so you know that the probablilty of rolling two is 1/6 right

OpenStudy (javalos01):

or am i wrong

OpenStudy (anonymous):

Yes, that is correct

OpenStudy (javalos01):

well it is confusing but what is your best guess

OpenStudy (javalos01):

we know all of them have a probibility of 1/6 so what u do with the 0.4 is divide or multiply

OpenStudy (anonymous):

Multiply?

OpenStudy (javalos01):

yes correct

OpenStudy (javalos01):

and the equation is

OpenStudy (javalos01):

0.4 *6 correct

OpenStudy (anonymous):

I thought you said multiply. So wouldn't it be 0.4 x 1/6?

OpenStudy (javalos01):

the * is the multivly sign

OpenStudy (javalos01):

and the / is division

OpenStudy (javalos01):

so after you multiply

OpenStudy (javalos01):

u get ??

OpenStudy (anonymous):

2.4?

OpenStudy (javalos01):

correct

OpenStudy (javalos01):

and thats your answer 2.4 you get it

OpenStudy (anonymous):

Where did the 6 come from though? I thought we were using 1/6. So wouldn't 0.4 x (1/6)= .0667?

OpenStudy (javalos01):

ik

OpenStudy (javalos01):

but we have to multiply th ewhole number and then

OpenStudy (anonymous):

I thought probability was supposed to be between 0 and 1

OpenStudy (javalos01):

@Rushwr

OpenStudy (javalos01):

no its one through 6

OpenStudy (javalos01):

so if yyou peice it all together 0.4x6=2.4 And that peices it all together

OpenStudy (javalos01):

u get it

OpenStudy (javalos01):

maybe you can help him understand rushwer

OpenStudy (rushwr):

Soo basically what u have to know is when u add all the probabilities they should be equal to 1

OpenStudy (javalos01):

yes they do 1/6+1/6 ect.

OpenStudy (javalos01):

and youll get 1 once you add them all and thank you so much rushwer for your help

OpenStudy (javalos01):

i really appreciate it

OpenStudy (rushwr):

P1 = probability of rolling 1 P2 = probability of rolling 2 P3 = probability of rolling 3 P4= probability of rolling 4 P5= probability of rolling 5 P6 = Probability of rolling 6 P1+P2+P3+P4+P5+P6 = 1

OpenStudy (javalos01):

yea thats what i meant too thats what i was trying to help him get in my equation

OpenStudy (rushwr):

Except for P6 others are same right? so we can say P1=P2=P3=P4=P5 so instead of P1+P2+P3+P4+P5 we can right 5P1 right?

OpenStudy (javalos01):

yea

OpenStudy (javalos01):

So do you get it now

OpenStudy (javalos01):

@tennis96

OpenStudy (rushwr):

so 5P1 + P6 = 1 We know P6 = 0.4 So 5P1 = 1-0.4 = 0.6 So P1 = 0.6/5 = 0.12 Except for the probability of rolling 6 other probabilities are 0.12

OpenStudy (javalos01):

yea correct

OpenStudy (anonymous):

But it says all other sides are equally likely. So that means the probability of rolling a 6 is 0.4. And the probability or rolling either a 1,2,3,4, or 5 is 0.4 as well?

OpenStudy (rushwr):

That is ; P1= 0.12 P2 = 0.12 P3 =0.12 P4 = 0.12 P5 = 0.12 P6 = 0.4 Now add up the probabilities of the even numbers ! That is P2+P4+P6 Then u get the probability of rolling an even number

OpenStudy (javalos01):

yea do you get it now tennis

OpenStudy (anonymous):

OHH it just clicked

OpenStudy (anonymous):

So all the other trials would have to be equal to 0.6 right? so 0.6/ 5= .12

OpenStudy (javalos01):

yes thank you so much for letting us help you

OpenStudy (javalos01):

i have a great time personaly helping students find there way

OpenStudy (anonymous):

Thank you!

OpenStudy (javalos01):

your wekcome thank you or letting us help you

OpenStudy (rushwr):

"The probability of rolling a 6 is 0.4 and all other sides are equally likely" this what the question said. I think u took it wrong !!! This means probability of rolling 6 is 0.4 and all the others equal. But this doesn't mean all the other sides are equal to 0.4 . Simply it says the probability of rolling 6 is different than the others

OpenStudy (anonymous):

Yeah I did haha! thanks for pointing this out to me Rusher!

OpenStudy (rushwr):

Rushwr* lol no problem !!!!

OpenStudy (javalos01):

can i have a medal plzz

OpenStudy (javalos01):

i helped too lol

OpenStudy (javalos01):

im just kidding i dont care about medals i just care about helping other students

OpenStudy (rushwr):

okai Done ! I gave u a medal !!!!!

OpenStudy (javalos01):

thnx

OpenStudy (rushwr):

I hope u got the answer right 11 @tennis96 Can I know ur answer before I leave this ?

OpenStudy (rushwr):

No problem :)

OpenStudy (javalos01):

yea thats what i asked too

OpenStudy (anonymous):

I'm guessing I add all the probabilities together. P(2)= .12 P(4)=.12 P(6)= .4. Added together is this ..

OpenStudy (anonymous):

.64?

OpenStudy (rushwr):

Correct !!!!

OpenStudy (anonymous):

Okay perfect thank you :)

OpenStudy (javalos01):

yes you are correct

OpenStudy (rushwr):

Anytym :)

OpenStudy (javalos01):

thank you for everything